OpenStudy (anonymous):

Help Please This Question Might Be Stupid. what is slope i know y=mx+b but how do you find slope and what does m and b stand for. i am really lost and i am 5 weeks behind math in FLVS.

OpenStudy (brooke2353):

boll umm

OpenStudy (anonymous):

If I remember correctly, you look at the y-intercept, and depending on the direction traveled, the slope can be positive or negative.

OpenStudy (anonymous):

If it's going bottom to top from left to right, it's positive. The inverse of Top to bottom from left to right is negative.

OpenStudy (brooke2353):

lets say y=3x+2 they slope would be 2 cause thats where the y inrcept is

OpenStudy (brooke2353):

it wont show

OpenStudy (anonymous):

you have to click on the link

OpenStudy (anonymous):

We don't have access to your course. You'll have to right click the image, download it, and link it via "Attack File".

OpenStudy (anonymous):

*attach lol

OpenStudy (anonymous):

its not in my coarse i clicked on it and it show it like a regular image

OpenStudy (brooke2353):

uhhh what?

OpenStudy (mathmale):

Your topic is "equations of straight lines." A typical equation for a straight line is y=mx+b. Stop a minute. Ask yourself (or look up the concept): What does "slope of the line" mean in this context? Next, ask yourself: which variable in y=mx+b represents the slope of the line? Next, substitute 0 for x in y=mx+b. After simplifying the result, what do you have? What does this y value signify?

OpenStudy (anonymous):

I mean we can't log in as you so we can't see the image directly from the page. :/

OpenStudy (anonymous):

its not on the page where my assignment it its like when you click on an image from googl e and click on the show full size kinda thing

OpenStudy (anonymous):

@mathmale i looked on google to see what slope of a line was and it only came up with y=mx+b

OpenStudy (brooke2353):

slope is the y

OpenStudy (brooke2353):

lol meant slope is the b

OpenStudy (anonymous):

ok so with this image what would the slope be 300?

OpenStudy (anonymous):

but what is b? b=slope ok but how do i find slope i swear its gonna be easy and im gonna look like a retard.

OpenStudy (brooke2353):

the slope is where the line touches the y intercept

OpenStudy (anonymous):

so 300? so y=b is what you said?

OpenStudy (anonymous):

@Aaran24 am i on the right track ?

OpenStudy (anonymous):

@AloneS can you help?

OpenStudy (anonymous):

It's been a few months since I've been doing this actually... working in Geometry. But I think you're making progress for sure.

OpenStudy (anonymous):

@Aaran24 but i dont understand it slope=300? and i put 300 where b is or m?

OpenStudy (anonymous):

Slope is m and the y-intercept is b

OpenStudy (anonymous):

I AM SO CONFUSED ok 1st slope was y then b now its m?

OpenStudy (anonymous):

y=300x+b ?

OpenStudy (anonymous):

Yes! Because m=slope and your slope is 300 !

OpenStudy (anonymous):

but y intercept is also 300 ?

OpenStudy (anonymous):

because it doesn't go through any other lines?

OpenStudy (anonymous):

How did you find your slope?

OpenStudy (anonymous):

someone said the slope was where the line touches the y so i said 300 they said yes

OpenStudy (anonymous):

this is what i wanted to know was how to find slope

OpenStudy (anonymous):

i did this last year in pre algebra but do i remember nooooope lol

OpenStudy (anonymous):

Slope is when you take two points and use the formula y2-y1/x2-x1

OpenStudy (anonymous):

so i would use 300 and 450?

OpenStudy (anonymous):

So chose any two points and use that formula

OpenStudy (anonymous):

300 2-300 1/ 450 2-450 1???

OpenStudy (anonymous):

thats confusing?

OpenStudy (anonymous):

I used 150 for x1 and 200 for y1 and then I used 300 for x2 and 100 for y2

OpenStudy (anonymous):

there is only 2 points did u click on the graph?

OpenStudy (anonymous):

@ILH_HIM_

OpenStudy (anonymous):

You should any points on that graph for example you can choose the points 150,200 and 300,200

OpenStudy (anonymous):

I mean 300,100

OpenStudy (anonymous):

So find those points on the graph and use the slope formula and you should get the answer

OpenStudy (anonymous):

oh ok thankss @ILH_HIM_